A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-06-04T07:29:50-0400

-7x + 3y = - 21.5

3y = 7x - 21.5

y = "\frac{7}{3}"x - "\frac{21.5}{3}"

Slope of line (m1) is 7/3

If the line is parallel than the Slope of the parallel line must be 7/3.

And

If the line is perpendicular than m1 × m2 = -1

"\frac{7}{3}" × m2 = -1

m2 = - "\frac{3}{7}"

Than the equation of line is

y = -"\frac{3}{7}" x - "\frac{21.5}{3}" .